# Dipak Meher    Carlotta Domeniconi

### v1

@@ V-B1 Node Duplication
Table I presents a comparison between CORE-KG, CoreKG-no-str-prompt, and the GraphRAG baseline. As shown in Table II, the removal of our structured prompt in CoreKG-no-str-prompt increases the node duplication rate from 20.27% (CORE-KG) to 21.15%, reflecting a relative degradation of approximately 4.3%. This highlights the role of CORE-KG’s structured prompt in guiding the LLM toward precise, domain-relevant entity extraction and in mitigating node duplication, even when coreference resolution is still present.
The structured prompt in CORE-KG employs sequential extraction and explicit type definitions to mitigate attention drift. This design prevents the LLM from attending to multiple entity and relationship types simultaneously, which can lead to over-extraction or conflated outputs. Without such structured guidance, as in CoreKG-no-str-prompt, the model may produce longer or imprecise spans that inflate the number of nodes. For example, in the sentence “The driver and the smuggler entered Texas with the group,” the model may extract relevant entities such as “the group” and “Texas,” but also over-extract the phrase “entered Texas with the group” as a separate entity. Although this phrase refers to the same underlying entities, its inclusion introduces duplication due to a loss of precision in entity boundaries, driven by attention drift.
Case-wise analysis further supports this conclusion. CoreKG-no-str-prompt shows higher duplication than CORE-KG in most cases. Notably, Case 06 demonstrates the largest increase, with node duplication rising from 6.25% in CORE-KG to 15.79% in CoreKG-no-str-prompt. This indicates that both structured prompting and coreference resolution are jointly necessary for achieving high-impact reductions in duplication.
@@ V-B2 Noisy Nodes
Figure 2: Comparison across CORE-KG, CoreKG-no-coref, and CoreKG-no-str-prompt graphs on a sample legal case. Duplicate entities are highlighted with rectangles, noisy entities with ovals, and disconnected nodes with dashed ovals.
@@ V-C Qualitative Analysis of Extracted Graphs
Figure 2 presents the generated graphs from CORE-KG, CoreKG-no-coref, and CoreKG-no-str-prompt for Case 6. This section analyzes the structural gains achieved through the incorporation of coreference resolution and structured prompts in the CORE-KG pipeline. For clarity, relationship labels are omitted in the visualizations, and node sizes are scaled based on their degree (i.e., the number of connected relationships).
@@ V-C1 Analysis of Relationship-to-Node Ratio
The graph generated by CoreKG-no-str-prompt contains the highest number of nodes and relationships, with 76 nodes and 127 edges. CoreKG-no-coref generates a moderately sized graph with 59 nodes and 107 edges, while CORE-KG produces a more compact structure containing 32 nodes and 92 edges. CORE-KG achieves the highest R/N ratio of 2.88, indicating a high degree of structural coherence. In contrast, CoreKG-no-coref and CoreKG-no-str-prompt yield lower ratios of 1.81 and 1.67, respectively. Although CoreKG-no-str-prompt produces more nodes, many appear weakly connected or isolated, suggesting increased noise. The higher density of the CORE-KG graph demonstrates the combined benefit of coreference resolution and structured prompting in producing a more informative and compact representation.
@@ V-C2 Per-Case Graph Comparison and Structural Gains
We further analyze the graph outputs by visually inspecting the impact of duplicate and noisy node generation. In Figure 2, duplicate nodes are marked using rectangles, while noisy nodes are enclosed in ovals. Duplicate nodes are colored identically to indicate that they represent the same entity but appear in multiple textual variations.
@@
For example, the main defendant in this case, “Richard Stone,” appears in several variations such as “Defendant Stone” and “Stone.” Similarly, other person-type entities like “Eleanor Grant,” “Victor Barnes,” and “Adam Everett” are each represented as multiple nodes due to their duplicate variations. Additionally, “John Ed Carlson,” listed as an “Undocumented Alien,” also appears in a shortened form as “Carlson,” further contributing to node duplication. This issue is not limited to person-type entities. Route-type entities such as “Interstate Highway 35” are inconsistently represented as “I-35” and “Interstate.” Location-type entities like “San Antonio” appear both as “San Antonio, Texas” and simply “San Antonio.” Likewise, means of transportation such as “Nissan Maxima” are scattered across forms like “Maxima,” “Car,” and “Vehicle.” These duplications—among many others—reduce clarity and introduce structural redundancy in the graph, emphasizing the importance of applying coreference resolution prior to knowledge graph construction.
Regarding noise, CoreKG-no-coref performs better than CoreKG-no-str-prompt and is nearly comparable to CORE-KG. The use of modified structured prompts in CoreKG-no-coref helps filter out irrelevant entities and ensures more precise entity and relationship extraction. However, some entities remain disconnected from the main graph, as shown in Figure 2(b). This is primarily due to the presence of multiple surface forms of the same entity in the text, which hinders proper merging during post-extraction alignment. Entities and relationships are typically merged based on matching names and types; however, when a location such as “Antonio” appears as a variation of “San Antonio,” the system fails to unify them, resulting in disconnected nodes.
Another contributing factor is attention drift caused by these multiple surface variations. Legal texts are inherently dense and contain numerous entities within short spans of text. The presence of variant forms dilutes the model’s attention, impacting the quality of entity and relationship extraction and increasing the likelihood of unconnected or missed nodes. For instance, the “65 mile marker” is properly linked to the main graph in CORE-KG (Figure 2(a)) and CoreKG-no-str-prompt (Figure 2(c)), but is absent in CoreKG-no-coref (Figure 2(b)).
@@
However, CoreKG-no-str-prompt suffered significantly from noisy entity extraction due to the use of GraphRAG’s standard prompt, which lacks the structured guidance implemented in CORE-KG. As shown in Figure 2(c), multiple government-related entities are unnecessarily extracted around nodes like Richard Stone” and Victor Barnes.” Additionally, several fine-grained but irrelevant entities appear, such as Window” and Car Frame,” both referring to parts of the “Nissan Maxima.” These cases illustrate the importance of in-prompt filtering, sequential extraction, and clearly defined entity types, all key components of CORE-KG’s structured prompt.
The standard GraphRAG prompt, by contrast, lacks structured mechanisms and therefore struggles with precise extraction. As previously discussed, this leads to issues such as attention drift, misallocation of focus, and noisy or incomplete outputs, especially in densely packed legal texts. These problems manifest in the CoreKG-no-str-prompt graph through unconnected and irrelevant nodes. CORE-KG’s structured prompt effectively addresses these challenges, resulting in cleaner and more coherent graphs.
@@ VI Conclusion
In this work, we conducted a systematic ablation study of CORE-KG to quantify the individual contributions of its two key components: type-aware coreference resolution and domain-guided structured prompts. By selectively disabling each module, we measured how their absence affects knowledge graph quality in terms of node duplication and legal noise. Removing coreference resolution increases node duplication by 28.32% and noisy nodes by 4.32%, while removing structured prompts increases node duplication by 4.34% and noisy nodes by 73.33%. These findings demonstrate that both modules are essential and complementary. Structured prompts play a dominant role in improving extraction precision, while coreference resolution enhances entity consolidation by addressing referential ambiguity. This clear separation in impact highlights the value of modular LLM pipelines that explicitly target different error modes. Future work may explore dynamic prompt adaptation, examine how cleaner KGs improve retrieval accuracy and answer coherence in downstream RAG systems, and investigate cross-case meta-coreference to construct unified super-graphs for tracking actors and patterns across legal documents.

### v2

@@ V-B1 Node Duplication
As shown in Table II, removing the structured prompt in CoreKG-no-str-prompt increases the node duplication rate from 20.28% (CORE-KG) to 21.15%, a relative degradation of approximately 4.29%. This highlights the role of CORE-KG’s structured prompt in guiding the LLM toward precise, domain-relevant entity extraction and in mitigating node duplication, even when coreference resolution is still present.
The structured prompt in CORE-KG employs sequential extraction and explicit type definitions to mitigate attention drift. This design prevents the LLM from attending to multiple entity and relationship types simultaneously, which can lead to over-extraction or conflated outputs. Without such structured guidance, as in CoreKG-no-str-prompt, the model may extract longer or imprecise entities that inflate the number of nodes. For example, in the sentence “The driver and the smuggler entered Texas with the group,” the model may extract relevant entities such as “the group” and “Texas,” but also over-extract the phrase “entered Texas with the group” as a separate entity. Although this phrase refers to the same underlying entities, its inclusion introduces duplication due to loss of precision in entity extraction, driven by attention drift.
Case-wise analysis (Figure 2) further supports this conclusion. CoreKG-no-str-prompt shows higher duplication than CORE-KG in most cases. Notably, Case 06 demonstrates the largest increase, with node duplication rising from 6.25% in CORE-KG to 15.79% in CoreKG-no-str-prompt. This suggests that structured prompting and coreference resolution are jointly necessary for achieving substantial reductions in duplication.
@@ V-B2 Noisy Nodes
Figure 2: Graphs of (a) CORE-KG; (b) CoreKG-no-coref; (c) CoreKG-no-struct-prompt, shown for Case 06. Duplicate entities are indicated with solid rectangles, noisy or irrelevant entities with solid ovals, and disconnected or weakly connected nodes with dashed ovals.
@@ V-C Qualitative Analysis of Extracted Graphs
Figure 2 presents the generated graphs from CORE-KG, CoreKG-no-coref, and CoreKG-no-str-prompt for Case 6. This section analyzes the structural gains achieved through the incorporation of coreference resolution and structured prompts in the CORE-KG pipeline. To provide a clean visual presentation of the graph, relationship labels are omitted in the visualizations, and node sizes are scaled based on their degree (i.e., the number of relationships connected to that node).
@@ V-C1 Analysis of Relationship-to-Node Ratio
The graph generated by CoreKG-no-str-prompt contains the largest number of nodes and relationships, with 76 nodes and 127 edges. CoreKG-no-coref produces a moderately sized graph with 59 nodes and 107 edges, while CORE-KG yields a more compact structure containing 32 nodes and 92 edges. CORE-KG achieves the highest R/N ratio of 2.88, indicating a high degree of structural coherence. In contrast, CoreKG-no-coref and CoreKG-no-str-prompt produce lower ratios of 1.81 and 1.67, respectively. Although CoreKG-no-str-prompt produces more nodes, many of them appear loosely connected or isolated, suggesting increased noise. The higher density of the CORE-KG graph demonstrates the combined benefit of coreference resolution and structured prompting in producing a more informative and compact representation.
@@ V-C2 Per-Case Graph Comparison and Structural Gains
We further analyze the graph outputs by visually inspecting the impact of duplicate and noisy node generation. In Figure 2, duplicate nodes are marked using rectangles, while noisy nodes are enclosed in ovals. Duplicate nodes are colored identically to indicate that they refer to the same real-world entity but appear under different surface forms in the text.
@@
For example, in the CoreKG-no-coref graph (Figure 2(b)), the main defendant, “Richard Stone,” appears in several surface variations, such as “Defendant Stone” and “Stone”. Similarly, other person-type entities, including “Eleanor Grant,” “Victor Barnes,” and “Adam Everett”, are each represented as multiple nodes due to these unresolved duplicate mentions. Additionally, “John Ed Carlson,” listed as an “Undocumented Alien,” also appears in a shortened form as “Carlson,” further contributing to node duplication. This issue is not limited to person-type entities. Route-type entities such as “Interstate Highway 35” are inconsistently represented as “I-35” and “Interstate.” Location-type entities like “San Antonio” appear both as “San Antonio, Texas” and simply “San Antonio.” Likewise, means of transportation such as “Nissan Maxima” are scattered across forms like “Maxima,” “Car,” and “Vehicle.” These duplications, among many others, reduce clarity and introduce structural redundancy in the graph, emphasizing the importance of applying coreference resolution prior to knowledge graph construction.
Regarding noise, CoreKG-no-coref performs better than CoreKG-no-struct-prompt and is nearly comparable to CORE-KG. The structured prompt still filters many irrelevant entities and supports more precise extraction. However, several entities remain disconnected from the main graph, as shown in Figure 2(b). This occurs for two reasons.
First, duplicate entities may appear in different surface forms, preventing correct merging. For example, the location “Antonio” is a shortened form of “San Antonio,” but since the merging step relies on matching written names, these are not unified and therefore appear as separate nodes. This observation highlights a limitation in the current entity merging component of the KG construction module, indicating that it could be made more robust.
Second, some nodes lack extracted relationships even when those relationships are present in the text. This occurs due to attention drift, where the model processes many entities in a dense sentence and overlooks certain links, leaving those nodes isolated in the graph. For example, the “65 mile marker” is properly linked to the main graph in CORE-KG (Figure 2(a)) and CoreKG-no-str-prompt (Figure 2(c)), but is missing in CoreKG-no-coref (Figure 2(b)). Structured prompting helps mitigate attention drift. As a result, CoreKG-no-coref contains around 4 isolated nodes, whereas CoreKG-no-str-prompt, which removes the structured prompt, shows approximately 12.
@@
However, CoreKG-no-str-prompt exhibits substantially higher noisy entity extraction, since it uses the standard GraphRAG prompt, which lacks the structured instructions that filter out irrelevant or noisy mentions. As shown in Figure 2(c), multiple government-related entities are unnecessarily extracted around nodes like “Richard Stone” and “Victor Barnes.” Additionally, several fine-grained but irrelevant entities appear, such as “Window” and “Car Frame,” both referring to parts of the “Nissan Maxima”. These cases illustrate the importance of in-prompt filtering, sequential extraction, and clearly defined entity types, all key components of CORE-KG’s structured prompt.
@@ VI Conclusion
In this work, we conducted a systematic ablation study of CORE-KG to quantify the individual contributions of its two key components: type-aware coreference resolution and domain-guided structured prompts. By isolating each module, we measured how its removal affects knowledge graph quality in terms of node duplication and legal noise. Removing coreference resolution increases node duplication by 28.25% and noisy nodes by 4.32%, while removing structured prompts increases node duplication by 4.29% and noisy nodes by 73.33%. These findings demonstrate that both modules are essential and complementary. Structured prompts play a dominant role in improving extraction precision, while coreference resolution enhances entity consolidation by resolving referential variation. This separation in impact highlights the value of modular LLM pipelines, where each component targets a distinct error mode. Future work may explore automating and optimizing prompt construction, evaluate how improved graph quality affects retrieval accuracy and answer coherence in downstream RAG systems, and develop cross-document coreference resolution to construct unified graphs that link recurring actors and patterns across legal cases.